OpenStudy (anonymous):

What is the solution to the equation 1/\[\sqrt{8}\] = 4(m – 3)

OpenStudy (anonymous):

\[1/\sqrt{8}=4(m-3)\]\[1/\sqrt{8}=4m-12\]\[1/\sqrt{8}+12=4m\]\[4/\sqrt{8}+3=m\]\[4\sqrt{8}/8+3=m\]\[\sqrt{8}/2+3=m\]\[4\sqrt{8}/8+3=m\]\[m=\sqrt{2}+3\]